An ice skater with moment of inertia 70.0 kg•m2 is spinning at 41.0 rpm. If the

skater pulls in her arms, her moment of inertia decreases to 50.0 kg•m2. What is

the skater’s resulting angular velocity?
Answer #1

apply angular momentum conservation

Li = Lf  

70*41 = 50 *w

w = 70*41/50 = 57.4 rpm

or 57.4 *2pi /60 = 6.01091394 rad/s

Goodluck for exams Comment in case any doubt, will reply for sure..